OpenStudy (anonymous):

find the surface area of the triangular prism. A- 81 in.^2 B- 70.8 in.^2 C- 63 in.^2 D- 78.6 in.^2

OpenStudy (anonymous):

its B

OpenStudy (anonymous):



OpenStudy (anonymous):

its not the best drawing sorry

OpenStudy (johnweldon1993):

Hmm seems good enough to me So surface area is going to be \[\large A = bh + 2LS + LB\] L = length B = base H = height S = side length So here we have \[\large A = (3)(2.6) + 2(7)(3) + (7)(3)\] \[\large A = 7.8 + 42 + 21\] \[\large A = 7.8 + 63\] \[\large A = 70.8\]

OpenStudy (anonymous):

thank you everyone!

OpenStudy (johnweldon1993):

anytime!